# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 3 Weight-Space Averaging and Geometric Interpolation Methods
This section examines weight-space averaging and geometric interpolation—the most fundamental class of merging techniques that operate directly on model parameters. We organize these approaches along a spectrum of geometric sophistication: static averaging with fixed coefficients, trajectory-based methods that exploit optimization dynamics, and geometric interpolation (e.g., SLERP) that respects the manifold structure of weight spaces.
@@ 3.1 Linear Averaging and Static Model Soups
Weight averaging represents the most fundamental instantiation of model merging. It computes the element-wise arithmetic mean of model parameters, treating weight space as a Euclidean vector space. This formula admits a natural generalization to weighted averaging $\theta_{\text{avg}}=\sum_{i}\alpha_{i}\theta_{i}$ with $\sum_{i}\alpha_{i}=1$, where the coefficients $\alpha_{i}$ can be tuned on a held-out set or determined via importance metrics such as Fisher information (Matena and Raffel, 2022). Given a collection of $N$ models with parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$, uniform weight averaging produces a merged model with parameters:
@@
This formulation assumes all source models share identical architectures with correspondingly positioned parameters. This works because of linear mode connectivity (Singh and Jaggi, 2020): models fine-tuned from a common pretrained initialization tend to reside within the same loss basin, enabling linear interpolation between solutions without traversing high-loss regions. Singh and Jaggi (2020) demonstrated that neural networks sharing early training trajectories exhibit this connectivity property, while (Izmailov et al., 2018) provided theoretical analysis connecting proximity to initialization with favorable generalization behavior. Importantly, this connectivity assumption breaks down when models originate from independent random initializations or when fine-tuning procedures cause parameters to diverge substantially from the shared pretrained manifold, resulting in interpolation paths that traverse high-loss barriers (Singh et al., 2024).
The Model Soups framework, introduced by Wortsman et al. (2022a), represents a principled advancement in weight-space averaging by systematically studying the aggregation of multiple fine-tuned checkpoints. The framework establishes two primary aggregation strategies: the uniform soup and the greedy soup (Jang et al., 2024).
The uniform soup computes the arithmetic mean of all $N$ available model checkpoints without selection:
@@
The greedy soup addresses this limitation through iterative conditional selection, incorporating models into the aggregate only when their inclusion improves held-out validation performance. Formally, starting with an initial model $\theta_{\text{soup}}=\theta_{1}$, each candidate $\theta_{i}$ is added only if:
@@
Empirically, Model Soups demonstrated substantial improvements across multiple benchmarks. While initially validated on vision models, this approach has since been applied to large language models. The key advantage is enhanced robustness to distribution shift. These improvements (Wortsman et al., 2022b) incur no additional inference cost compared to deploying a single model, as the merged architecture maintains identical computational requirements during deployment.
This selective aggregation strategy addresses a fundamental limitation of uniform averaging: not all models contribute equally to the merged solution, and certain checkpoints may introduce parameter interference that degrades performance. Wortsman et al. (2022a) demonstrate that weight-space operations can enable constructive task composition when models are appropriately combined, while unstructured aggregation frequently results in destructive interference patterns. Model Ratatouille (Ram’e et al., 2023a) extends this insight by demonstrating that recycling diverse auxiliary models fine-tuned on various domains improves out-of-distribution generalization, providing a principled approach for leveraging the growing ecosystem of available fine-tuned models. The relationship between model aggregation and federated learning has been explored extensively, with Li et al. (2023a) examining how weighted averaging strategies can be adapted to scenarios involving heterogeneous client contributions (Li et al., 2020; Wang et al., 2020) and non-identically distributed data.
@@
Unlike exhaustive combinatorial search over model subsets, which becomes computationally prohibitive as the candidate pool grows exponentially, greedy selection provides an efficient approximation that scales linearly with the number of available checkpoints. However, this approach assumes that validation performance serves as a reliable proxy for generalization, an assumption that may not hold under significant distribution shift between validation and deployment domains. Furthermore, theoretical analyses indicate that linear averaging methods, while computationally efficient, do not account for the permutation symmetries inherent in neural network weight spaces, motivating the development of alignment-based interpolation techniques discussed in later subsections. Beyond these structural considerations, the uniform weighting scheme inherent in these approaches raises a more fundamental question: whether treating all parameters as equally important accurately reflects their heterogeneous contributions to model functionality.
@@ 3.2 Importance-Weighted Averaging and Fisher Information
Importance-weighted averaging methods directly address this question by rejecting the implicit assumption that all parameters contribute equally to model functionality, an assumption that significantly misrepresents the heterogeneous nature of neural network weight spaces. In practice, certain parameters encode critical task-specific information while others remain largely unchanged from pretraining or exhibit high variance across fine-tuning runs. By incorporating second-order information about parameter sensitivity, these methods enable more principled aggregation that prioritizes high-information parameters while attenuating noise from less consequential weights.
@@
The theoretical foundation for importance-weighted merging derives from the Fisher Information Matrix (FIM), which characterizes the local curvature of the loss landscape regarding model parameters. For a model with parameters $\theta$ and loss function $\mathcal{L}$, the Fisher Information Matrix is defined as $F=\mathbb{E}[\nabla_{\theta}\log p(x|\theta)\nabla_{\theta}\log p(x|\theta)^{\top}]$, where the expectation is taken over the data distribution. Parameters associated with high Fisher information values correspond to directions in weight space where small perturbations significantly affect model predictions, indicating their importance for task performance. Fisher-weighted averaging exploits this insight by computing the merged parameters as $\theta_{\text{merged}}=(\sum_{i}F_{i})^{-1}\sum_{i}F_{i}\theta_{i}$, where $F_{i}$ denotes the Fisher Information Matrix computed for model $i$. This formulation effectively weights each model’s contribution inversely proportional to its uncertainty, yielding a maximum a posteriori estimate under Gaussian assumptions about parameter distributions.
@@
The computational demands of full Fisher Information Matrix estimation render it intractable for contemporary large language models with billions of parameters. Practical implementations therefore employ diagonal approximations, where only the diagonal entries $F_{i,jj}$ are retained, reducing both memory requirements and computational complexity from $O(d^{2})$ to $O(d)$ where $d$ represents the parameter count. This diagonal Fisher-weighted averaging computes element-wise weighted means: $\theta_{\text{merged},j}=(\sum_{i}F_{i,jj}\theta_{i,j})/(\sum_{i}F_{i,jj})$, treating each parameter independently. While this approximation discards correlation structure between parameters, empirical investigations demonstrate that diagonal Fisher weighting substantially outperforms uniform averaging across diverse task combinations (Matena and Raffel, 2022).
@@
Alternative formulations leverage covariance statistics rather than Fisher information to estimate parameter importance. The RegMean approach (Jin et al., 2023) computes inner products between model activations (Matena and Raffel, 2022) to construct weight matrices that account for the representational structure learned by each model. Specifically, RegMean solves a linear regression problem that minimizes the expected squared difference between the merged model’s outputs and those of the constituent models, weighted by activation covariances. This representation-aware weighting implicitly captures parameter dependencies that diagonal Fisher approximations neglect, potentially yielding superior merging outcomes when models have learned complementary feature representations. WiSE-FT (Wortsman et al., 2022b) demonstrates that robust fine-tuning of zero-shot models can be achieved through simple weight interpolation between the pretrained and fine-tuned parameters, providing a principled approach for preserving zero-shot capabilities while acquiring task-specific expertise. Bayesian perspectives on weighted aggregation naturally extend these importance-based schemes, treating merged parameters as posterior estimates given observations from multiple training procedures. Recent work has also explored the relationship between model merging and continual learning (He et al., 2024; Jordan et al., 2022), with Stoica et al. (2023) proposing ZipIt! for merging models trained on disjoint tasks by exploiting feature correlations, and Tang et al. (2024a) providing systematic benchmarking of merging methods under controlled conditions.
@@
The relationship between model merging and catastrophic forgetting (Kirkpatrick et al., 2016) presents another critical consideration. When fine-tuned models are merged, the resulting model must retain capabilities from all constituent tasks without catastrophically forgetting any individual specialization. Elastic Weight Consolidation (EWC) and related continual learning methods provide theoretical frameworks for understanding which parameters are most critical for specific capabilities, insights that directly inform importance-weighted merging strategies.
@@
Despite their theoretical appeal, importance-weighted methods introduce several practical challenges. Fisher information estimates require access to representative data samples from each constituent model’s training distribution, information that may be unavailable when merging pretrained models obtained from external sources. Furthermore, the approximation quality of diagonal Fisher estimates degrades in highly overparameterized networks where off-diagonal correlations capture essential structural relationships (Matena and Raffel, 2022). Unlike uniform averaging, which requires only the model weights themselves, importance-weighted approaches demand additional computational overhead for Fisher or covariance estimation, creating tradeoffs between merging quality and practical deployability that practitioners must navigate based on their specific resource constraints and accuracy requirements. Crucially, both uniform and importance-weighted averaging share a common assumption: they operate on models trained independently, treating the aggregation problem as combining discrete endpoints without exploiting any structure from the optimization process itself.
@@
Despite its simplicity and empirical success, linear averaging exhibits well-characterized failure modes. When source models are fine-tuned with substantially different learning rates or regularization strengths, the resulting parameter vectors may reside in distinct loss basins, causing their average to land on a high-loss barrier (Frankle et al., 2020). Performance also degrades when the number of merged models grows large and the tasks are semantically diverse, a phenomenon termed task interference. Ramé et al. (Ram’e et al., 2023a) partially address this by recycling intermediate fine-tuned checkpoints from diverse auxiliary tasks (Model Ratatouille), demonstrating that averaging models exposed to a broader training distribution yields more robust merged solutions. Nevertheless, characterizing the precise conditions under which linear averaging provably succeeds remains an open theoretical challenge (see Section 2).
@@ 3.3 Trajectory-Based Averaging and Optimization Dynamics
Trajectory-based averaging represents a significantly distinct paradigm that directly exploits the temporal structure of optimization, precisely the information that endpoint-focused methods discard. Rather than combining final checkpoints of separate training procedures, these techniques aggregate intermediate states collected along the optimization trajectory, leveraging the insight that weight-space averages of sequential iterates can approximate solutions residing near the center of flat loss basins. This geometric centering yields models exhibiting enhanced robustness to distribution shifts and improved generalization characteristics.
Polyak Averaging. The theoretical foundations for trajectory-based averaging trace to Polyak and Juditsky’s seminal work on iterate averaging for stochastic approximation. Their analysis established that averaging SGD iterates achieves optimal asymptotic convergence rates, with the averaged solution $\bar{\theta}_{n}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}$ converging at rate $O(1/n)$ compared to $O(1/\sqrt{n})$ for the final iterate alone. This classical result provides the theoretical underpinning for modern trajectory averaging methods, demonstrating that temporal averaging serves as an implicit regularizer that suppresses gradient noise while preserving convergence to stationary points.
@@
SWAG and Uncertainty Quantification. Stochastic Weight Averaging Gaussian (SWAG) extends SWA to approximate Bayesian inference by additionally capturing the covariance structure of the optimization trajectory. Beyond computing the mean $\theta_{\text{SWA}}$, SWAG maintains a low-rank approximation to the covariance matrix $\Sigma\approx\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{i}-\theta_{\text{SWA}})(\theta_{i}-\theta_{\text{SWA}})^{\top}$, enabling posterior sampling for uncertainty estimation. This extension proves particularly valuable for calibrated predictions and out-of-distribution detection, achieving competitive performance with ensemble methods at substantially reduced computational cost.
Lookahead Optimizer. The Lookahead optimizer provides a complementary perspective by integrating trajectory averaging directly into the optimization procedure rather than applying it post-hoc. Lookahead maintains slow weights $\phi$ updated every $k$ steps toward the fast weights $\theta$ via $\phi\leftarrow\phi+\alpha(\theta-\phi)$, effectively performing online trajectory averaging. This approach reduces variance in gradient estimates and demonstrates improved stability across learning rate choices, often achieving faster convergence while maintaining or improving final performance.
@@
Applications in LLM Training. Trajectory averaging has demonstrated substantial benefits in large-scale language model development. In neural machine translation and general LLM pretraining, checkpoint averaging of the final training checkpoints is commonly used (Izmailov et al., 2018). In the model merging context, averaging multiple checkpoints from the same fine-tuning run can improve generation quality and robustness on downstream NLP tasks (Wortsman et al., 2022a).
The connection between trajectory-based averaging and broader model merging frameworks emerges when considering multiple training runs initialized from common starting points. When models share a pretrained initialization and undergo fine-tuning on related tasks, their optimization trajectories remain confined to overlapping basins of attraction, enabling checkpoints to be meaningfully averaged despite originating from nominally independent procedures. This perspective bridges trajectory-based methods with task vector approaches, as temporal averaging within a single fine-tuning run and spatial averaging across related runs both operate by centering solutions within shared loss basins, with combined approaches yielding additive benefits.
Practical implementations of trajectory-based averaging require attention to checkpoint selection strategy and computational overhead. Regarding collection frequency, Izmailov et al. established that averaging every epoch during the final 25% of training provides near-optimal results, while more frequent collection yields diminishing returns (Izmailov et al., 2018). For memory efficiency, running average formulations avoid storing multiple checkpoints: SWA requires only two model copies (current and averaged), while EMA maintains constant memory overhead regardless of training duration. The robustness benefits to distribution shift have been validated in multiple studies, with SWA-averaged models exhibiting smaller performance degradation on domain-shifted evaluation sets compared to single checkpoints, attributed to the implicit regularization toward flatter minima.
Despite their effectiveness, trajectory-based methods share a common characteristic with the importance-weighted approaches discussed earlier: they perform arithmetic averaging in weight space, implicitly treating the parameter manifold as Euclidean. This assumption, while reasonable for iterates within a single optimization trajectory that remain geometrically proximate, becomes increasingly problematic when averaging models that have diverged substantially during independent fine-tuning, a limitation that motivates geometric interpolation techniques designed to preserve the statistical properties of constituent parameter sets.
@@
Trajectory-based methods significantly require access to intermediate checkpoints from a single training run, making them inapplicable when only final model weights are available—the common scenario in open-weight model repositories. Furthermore, the memory overhead of storing $T$ checkpoints ($T\times|\theta|$ parameters) can be prohibitive for billion-parameter models, though rank-1 covariance approximations such as those in SWAG partially mitigate this.

### 30 Mar 2026 v2

@@ 3 Weight-Space Averaging and Geometric Interpolation Methods
This section examines weight-space averaging and geometric interpolation, the most fundamental class of merging techniques. We organize these approaches along a spectrum of geometric sophistication, from static averaging with fixed coefficients, through trajectory-based methods that exploit optimization dynamics, to geometric interpolation (e.g., SLERP) that respects the manifold structure of weight spaces.
@@ 3.1 Linear Averaging and Static Model Soups
Weight averaging represents the most fundamental instantiation of model merging. It computes the element-wise arithmetic mean of model parameters, treating weight space as a Euclidean vector space. This formula admits a natural generalization to weighted averaging $\theta_{\text{avg}}=\sum_{i}\alpha_{i}\theta_{i}$ with $\sum_{i}\alpha_{i}=1$, where the coefficients $\alpha_{i}$ can be tuned on a held-out set or determined via importance metrics such as Fisher information (Matena and Raffel, 2022). Given a collection of NN models with parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$, uniform weight averaging produces a merged model with parameters:
@@
This formulation assumes all source models share identical architectures with correspondingly positioned parameters. This works because of linear mode connectivity (Frankle et al., 2020), whereby models fine-tuned from a common pretrained initialization tend to reside within the same loss basin, so that linear interpolation between solutions avoids high-loss regions. However, this connectivity assumption breaks down when models originate from independent random initializations or when fine-tuning procedures cause parameters to diverge substantially from the shared pretrained manifold, resulting in interpolation paths that traverse high-loss barriers (Singh et al., 2024). For models lacking shared initialization, optimal transport methods (Singh and Jaggi, 2020) can align hidden units before averaging.
The Model Soups framework, introduced by Wortsman et al. (2022a), advances weight-space averaging by systematically studying the aggregation of multiple fine-tuned checkpoints. The framework establishes two primary aggregation strategies, the uniform soup and the greedy soup.
The uniform soup computes the arithmetic mean of all NN available model checkpoints without selection.
@@
The greedy soup overcomes this limitation through iterative conditional selection, incorporating models into the aggregate only when their inclusion improves held-out validation performance. Formally, starting with an initial model $\theta_{\text{soup}}=\theta_{1}$, each candidate $\theta_{i}$ is added only if
@@
Empirically, greedy soup consistently outperformed both uniform soup and individual fine-tuned models across multiple benchmarks, including ImageNet and its distribution-shifted variants. While initially validated on vision models, this approach has since been applied to large language models. The key advantage is enhanced robustness to distribution shift. These improvements (Wortsman et al., 2022a) incur no additional inference cost compared to deploying a single model, as the merged architecture maintains identical computational requirements during deployment.
This selective aggregation strategy addresses a fundamental limitation of uniform averaging, namely that not all models contribute equally, and certain checkpoints introduce parameter interference that degrades performance. Wortsman et al. (2022a) show that weight-space operations can enable constructive task composition when models are appropriately combined, while unstructured aggregation frequently results in destructive interference. Several extensions have broadened the applicability of this framework. ColD Fusion (Don-Yehiya et al., 2023) adapts the approach to distributed multi-task fine-tuning, showing that iterative collaborative descent through weight averaging supports continuous base model improvement without data sharing across participants. Model Ratatouille (Ram’e et al., 2023a) recycles diverse auxiliary models as initializations for parallel fine-tunings on the target task before averaging, improving out-of-distribution generalization. Transformer Fusion with Optimal Transport (Imfeld et al., 2024) uses optimal transport for neuron matching before averaging to handle alignment challenges specific to Transformer architectures.
Beyond these core extensions, multiple lines of work expand the frontier of weight-space merging. Bayesian approaches to checkpoint selection (Liu et al., 2024a) maximize intermediate checkpoint value during LLM pretraining through careful aggregation. Soup-of-Experts (Ablin et al., 2025) learns to linearly combine a bank of expert parameters, enabling test-time instantiation of specialist models without retraining, while Extrapolation Merging (Lin et al., 2025) shows that performance can continue improving beyond simple interpolation through extrapolation in weight space. Researchers have also explored the relationship between model aggregation and federated learning extensively, adapting weighted averaging strategies to scenarios involving heterogeneous client contributions (Li et al., 2020; Wang et al., 2020) and non-identically distributed data.
@@
Unlike exhaustive combinatorial search over model subsets, which becomes computationally prohibitive as the candidate pool grows exponentially, greedy selection provides an efficient approximation that scales linearly with the number of available checkpoints. However, this approach assumes that validation performance serves as a reliable proxy for generalization, an assumption that may not hold under substantial distribution shift between validation and deployment domains. Furthermore, theoretical analyses indicate that linear averaging methods, while computationally efficient, do not account for the permutation symmetries inherent in neural network weight spaces, motivating the development of alignment-based interpolation techniques discussed in later subsections. Beyond these structural considerations, the uniform weighting scheme treats all parameters as equally important, ignoring their heterogeneous contributions to model functionality, a limitation that importance-weighted approaches directly address.
@@ 3.2 Importance-Weighted Averaging and Fisher Information
Importance-weighted averaging methods confront this limitation directly by weighting each parameter according to its estimated contribution to model performance. In practice, certain parameters encode critical task-specific information while others remain largely unchanged from pretraining or exhibit high variance across fine-tuning runs. By incorporating second-order information about parameter sensitivity, these methods prioritize high-information parameters while attenuating noise from less consequential weights.
@@
Fisher-weighted merging draws its theoretical foundation from the Fisher Information Matrix (FIM), which characterizes the local curvature of the loss landscape regarding model parameters. For a model with parameters $\theta$ and loss function $\mathcal{L}$, the Fisher Information Matrix is defined as $F=\mathbb{E}[\nabla_{\theta}\log p(x|\theta)\nabla_{\theta}\log p(x|\theta)^{\top}]$, where the expectation is taken over the data distribution. Parameters associated with high Fisher information values correspond to directions in weight space where small perturbations strongly affect model predictions, indicating their importance for task performance. Fisher-weighted averaging exploits this insight by computing the merged parameters as $\theta_{\text{merged}}=(\sum_{i}F_{i})^{-1}\sum_{i}F_{i}\theta_{i}$, where $F_{i}$ denotes the Fisher Information Matrix computed for model ii. This formulation effectively weights each model’s contribution inversely proportional to its uncertainty, yielding a maximum a posteriori estimate under Gaussian assumptions about parameter distributions.
@@
Full Fisher Information Matrix estimation imposes computational demands that render it intractable for contemporary large language models with billions of parameters. Practical implementations therefore use diagonal approximations, where only the diagonal entries $F_{i,jj}$ are retained, reducing both memory requirements and computational complexity from $O(d^{2})$ to $O(d)$ where $d$ represents the parameter count. This diagonal Fisher-weighted averaging computes element-wise weighted means: $\theta_{\text{merged},j}=(\sum_{i}F_{i,jj}\theta_{i,j})/(\sum_{i}F_{i,jj})$, treating each parameter independently. While this approximation discards correlation structure between parameters, empirical investigations show that diagonal Fisher weighting consistently outperforms uniform averaging across diverse task combinations (Matena and Raffel, 2022).
@@
Alternative formulations use covariance statistics rather than Fisher information to estimate parameter importance. The RegMean approach (Jin et al., 2023) computes inner products between model activations to construct weight matrices that account for the representational structure learned by each model. Specifically, RegMean solves a linear regression problem that minimizes the expected squared difference between the merged model’s outputs and those of the constituent models, weighted by activation covariances. This representation-aware weighting implicitly captures parameter dependencies that diagonal Fisher approximations neglect, potentially yielding superior merging outcomes when models have learned complementary feature representations. WiSE-FT (Wortsman et al., 2022b) shows that robust fine-tuning of zero-shot models can be obtained through simple weight interpolation between the pretrained and fine-tuned parameters, providing a rigorous approach for preserving zero-shot capabilities while acquiring task-specific expertise. Bayesian perspectives on weighted aggregation naturally extend these importance-based schemes, treating merged parameters as posterior estimates given observations from multiple training procedures. Localize-and-Stitch (He et al., 2025a) introduces localized sparse merging, identifying tiny regions (1% of parameters) essential for each task and stitching only those back into the pretrained model, enabling flexible and continual skill composition with minimal storage overhead. Stoica et al. (2024) propose ZipIt! for merging models trained on disjoint tasks by exploiting feature correlations, and Yadav et al. (2025b) provide systematic benchmarking of merging methods under controlled conditions.
@@
Catastrophic forgetting (Kirkpatrick et al., 2017) presents another critical consideration for model merging. When fine-tuned models are merged, the resulting model must retain capabilities from all constituent tasks without catastrophically forgetting any individual specialization. Fuse to Forget (Zaman et al., 2024) shows that model merging can be strategically used for bias reduction and selective memorization, revealing that the forgetting properties of merging can be harnessed as a feature rather than a limitation. Merging by Matching Models in Task Parameter Subspaces (Tam et al., 2024a) proposes identifying and aligning task-specific parameter subspaces before merging, yielding improved capability preservation. Elastic Weight Consolidation (EWC) and related continual learning methods provide theoretical frameworks for understanding which parameters are most critical for specific capabilities, insights that directly inform importance-weighted merging strategies.
@@
Despite their theoretical appeal, importance-weighted methods introduce several practical challenges. Fisher information estimates require access to representative data samples from each constituent model’s training distribution, information that may be unavailable when merging pretrained models obtained from external sources. Moreover, the approximation quality of diagonal Fisher estimates degrades in highly overparameterized networks where off-diagonal correlations capture essential structural relationships (Matena and Raffel, 2022). Unlike uniform averaging, which requires only the model weights themselves, importance-weighted approaches demand additional computational overhead for Fisher or covariance estimation, creating tradeoffs between merging quality and practical deployability that practitioners must navigate based on their specific resource constraints and accuracy requirements. Crucially, both uniform and importance-weighted averaging share a common assumption in that they operate on models trained independently, treating the aggregation problem as combining discrete endpoints without exploiting any structure from the optimization process itself.
@@
Despite its simplicity and empirical success, linear averaging exhibits well-characterized failure modes. When source models are fine-tuned with markedly different learning rates or regularization strengths, the resulting parameter vectors may reside in distinct loss basins, causing their average to land on a high-loss barrier (Frankle et al., 2020). Performance also degrades when the number of merged models grows large and the tasks are semantically diverse, a phenomenon termed task interference. Ram’e et al. (2023a) partially address this by recycling intermediate fine-tuned checkpoints from diverse auxiliary tasks (Model Ratatouille), showing that averaging models exposed to a broader training distribution yields more robust merged solutions. At larger scales, Khalifa et al. (2024) show that optimizing merging weights across a pool of checkpoints at the 100B-parameter scale can recover near-Pareto-optimal performance by recycling suboptimal checkpoints that would otherwise be discarded. Nevertheless, characterizing the precise conditions under which linear averaging provably succeeds remains an open theoretical challenge (see Section 2).
@@ 3.3 Trajectory-Based Averaging and Optimization Dynamics
Trajectory-based averaging represents a fundamentally distinct approach that directly exploits the temporal structure of optimization, precisely the information that endpoint-focused methods discard. Rather than combining final checkpoints of separate training procedures, these techniques aggregate intermediate states collected along the optimization trajectory, exploiting the insight that weight-space averages of sequential iterates can approximate solutions residing near the center of flat loss basins. This geometric centering yields models exhibiting enhanced robustness to distribution shifts and improved generalization characteristics.
Polyak Averaging. The theoretical foundations for trajectory-based averaging trace to Polyak and Juditsky’s seminal work on iterate averaging for stochastic approximation. Their analysis established that averaging SGD iterates achieves optimal asymptotic convergence rates, with the averaged solution $\bar{\theta}_{n}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}$ converging at rate $O(1/n)$ for strongly convex objectives, compared to $O(1/\sqrt{n})$ for the final iterate alone. This classical result provides the theoretical underpinning for modern trajectory averaging methods, showing that temporal averaging serves as an implicit regularizer that suppresses gradient noise while preserving convergence to stationary points.
@@
SWAG and Uncertainty Quantification. Stochastic Weight Averaging Gaussian (SWAG) extends SWA to approximate Bayesian inference by additionally capturing the covariance structure of the optimization trajectory. Beyond computing the mean $\theta_{\text{SWA}}$, SWAG maintains a low-rank approximation to the covariance matrix $\Sigma\approx\frac{1}{n-1}\sum_{i=1}^{n}(\theta_{i}-\theta_{\text{SWA}})(\theta_{i}-\theta_{\text{SWA}})^{\top}$, enabling posterior sampling for uncertainty estimation. This extension proves particularly valuable for calibrated predictions and out-of-distribution detection, achieving competitive performance with ensemble methods at considerably reduced computational cost.
Lookahead Optimizer. The Lookahead optimizer provides a complementary perspective by integrating trajectory averaging directly into the optimization procedure rather than applying it post-hoc. Lookahead maintains slow weights $\phi$ updated every $k$ steps toward the fast weights $\theta$ via $\phi\leftarrow\phi+\alpha(\theta-\phi)$, effectively performing online trajectory averaging. This approach reduces variance in gradient estimates and shows improved stability across learning rate choices, often reaching faster convergence while maintaining or improving final performance.
@@
Applications in LLM Training. Trajectory averaging has shown clear benefits in large-scale language model development. In neural machine translation, checkpoint averaging of the final training checkpoints is standard practice (Vaswani et al., 2017), and SWA has been successfully applied to improve generalization in image classification (Izmailov et al., 2018). In the model merging context, averaging models from different fine-tuning runs with varying hyperparameters can improve robustness on downstream tasks (Wortsman et al., 2022a).
A bridge between trajectory-based averaging and broader model merging frameworks emerges when considering multiple training runs initialized from common starting points. When models share a pretrained initialization and undergo fine-tuning on related tasks, their optimization trajectories remain confined to overlapping basins of attraction, so that checkpoints can be meaningfully averaged despite originating from nominally independent procedures. This perspective bridges trajectory-based methods with task vector approaches, as temporal averaging within a single fine-tuning run and spatial averaging across related runs both operate by centering solutions within shared loss basins, with combined approaches yielding additive benefits.
Practical implementations of trajectory-based averaging require attention to checkpoint selection strategy and computational overhead. Regarding collection frequency, Izmailov et al. (2018) established that averaging checkpoints collected during the later stages of training with cyclic or constant learning rates provides near-optimal results, while more frequent collection yields diminishing returns. For memory efficiency, running average formulations avoid storing multiple checkpoints. SWA requires only two model copies (current and averaged), while EMA maintains constant memory overhead regardless of training duration. The robustness benefits to distribution shift have been validated in multiple studies, with SWA-averaged models exhibiting smaller performance degradation on domain-shifted evaluation sets compared to single checkpoints, attributed to the implicit regularization toward flatter minima.
Despite their effectiveness, trajectory-based methods share a common characteristic with the importance-weighted approaches discussed earlier, in that they perform arithmetic averaging in weight space, implicitly treating the parameter manifold as Euclidean. This assumption, while reasonable for iterates within a single optimization trajectory that remain geometrically proximate, becomes increasingly problematic when averaging models that have diverged substantially during independent fine-tuning, a limitation that motivates geometric interpolation techniques designed to preserve the statistical properties of constituent parameter sets.
@@
Trajectory-based methods inherently require access to intermediate checkpoints from a single training run, making them inapplicable when only final model weights are available, as is typical in open-weight model repositories. In addition, the memory overhead of storing $T$ checkpoints ($T\times|\theta|$ parameters) can be prohibitive for billion-parameter models, though rank-1 covariance approximations such as those in SWAG partially mitigate this.